# Downlink and Uplink ISAC in Continuous-Aperture Array (CAPA) Systems

### v1

@@ III-D Pareto-Optimal Design
In practical applications, the beamformer $w({\mathbf{t}})$ can be designed to meet various quality-of-service (QoS) requirements, which creates a trade-off between communication and sensing performance. To evaluate this trade-off, we investigate the Pareto boundary of the achievable SR-CR region in the downlink CAPA-based ISAC system. The Pareto boundary comprises SR-CR pairs where any improvement in one rate can only be achieved by sacrificing performance in the other. This boundary characterizes the optimal trade-off between communication and sensing capabilities, and provides valuable insights for performance optimization.
It is worth noting that the donwlink CR is a monotone increasing function with respect to $\hat{\gamma}_{\rm{c}}({{w}}({\mathbf{t}}))=\lvert\int_{{\mathcal{A}}_{\rm{t}}}h_{\rm{d}}({\mathbf{t}}){{w}}({\mathbf{t}}){\rm{d}}{\mathbf{t}}\rvert^{2}$, while the SR is a monotone increasing function with respect to $\hat{\gamma}_{\rm{s}}({{w}}({\mathbf{t}}))=\lvert\int_{{\mathcal{A}}_{\rm{t}}}a_{\rm{t}}(\mathbf{t})w({\mathbf{t}}){\rm{d}}{\mathbf{t}}\rvert^{2}$. Therefore, the achievable SR-CR region can be determined from the region defined by $\hat{\gamma}_{\rm{c}}({{w}}({\mathbf{t}}))$ and $\hat{\gamma}_{\rm{s}}({{w}}({\mathbf{t}}))$. In particular, any pair located on the Pareto boundary of this region can be identified by solving the following problem [40]:
@@
Please refer to Appendix -B for more details. ∎
@@
Although problem (58) is more mathematically tractable than problem (54), it remains non-convex. To solve this, we derive a closed-form solution using the Karush–Kuhn–Tucker (KKT) conditions [41].
@@
Please refer to Appendix -C for more details. ∎
@@ IV-A SIC-Based Receive Processing
$\begin{split}y_{\ell}({\mathbf{r}})&\approx\sqrt{P_{\rm{s}}}s_{{\rm{s}},\ell}a_{\rm{r}}(\mathbf{r})\beta_{\rm{s}}\int_{{\mathcal{A}}_{\rm{t}}}a_{\rm{t}}(\mathbf{t})w({\mathbf{t}}){\rm{d}}{\mathbf{t}}\\ &+\sqrt{P_{\rm{c}}}s_{{\rm{c}},\ell}\lvert{\mathcal{A}}\rvert h_{\rm{u}}(\mathbf{r})w_{\rm{u}}(\mathbf{p}_{\rm{c}})+n_{\ell}(\mathbf{r}),\end{split}$ (63)
where $\int_{\mathcal{A}}\lvert w_{\rm{u}}(\mathbf{u})\rvert^{2}{\rm{d}}{\mathbf{u}}=1\approx\lvert w_{\rm{u}}(\mathbf{p}_{\rm{c}})\rvert^{2}\lvert{\mathcal{A}}\rvert$ and $h_{\rm{u}}({\mathbf{r}})\triangleq h_{\rm{c}}({\mathbf{r}},\mathbf{p}_{\rm{c}})$. Without loss of generality, we set $w_{\rm{u}}(\mathbf{u})=w_{\rm{u}}(\mathbf{p}_{\rm{c}})=\frac{1}{\sqrt{\lvert{\mathcal{A}}\rvert}}$, which further simplifies $y_{\ell}({\mathbf{r}})$ to the follows:
@@ IV-B1 Performance of Sensing
From a worst-case design perspective, the following aggregate interference-plus-noise term:
@@
is treated as Gaussian noise [43], with $z_{{\rm{c}},\ell}\sim{\mathcal{CN}}(0,\sigma_{{\rm{c}},\ell}^{2})$ and
@@
$a_{\rm{r}}(\mathbf{r})$ $=a_{\rm{r},1}\psi_{1}(\mathbf{r})+a_{\rm{r},2}\psi_{2}(\mathbf{t}).$ (70c)
@@
where ${\mathbf{v}}=[v_{1},v_{2}]^{\mathsf{T}}$, ${\mathbf{h}}_{\rm{u}}=[h_{\rm{u},1},h_{\rm{u},2}]^{\mathsf{T}}$, and ${\mathbf{a}}_{\rm{r}}=[a_{\rm{r},1},a_{\rm{r},2}]^{\mathsf{T}}$. This is now a classical matrix-form maximization of a generalized Rayleigh quotient, whose optimal solution is given by [44]
@@
Using this solution, we derive the optimal value of the objective function as follows.
@@
The maximum value of $\frac{\lvert\int_{{\mathcal{A}}_{\rm{r}}}v_{\rm{s}}^{*}({\mathbf{r}})a_{\rm{r}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{\sigma_{{\rm{c}},\ell}^{2}}$ equals
$\frac{1}{\sigma^{2}}\left(\int_{\mathcal{A}_{\mathrm{r}}}{\!}\lvert a_{\mathrm{r}}(\mathbf{r})\rvert^{2}\mathrm{d}\mathbf{r}-\frac{\tilde{\gamma}_{\mathrm{c}}\lvert\int_{\mathcal{A}_{\mathrm{r}}}{a_{\mathrm{r}}(\mathbf{r})h_{\mathrm{u}}^{*}(\mathbf{r})}\mathrm{d}\mathbf{t}\rvert^{2}}{1+\tilde{\gamma}_{\mathrm{c}}\int_{\mathcal{A}_{\mathrm{r}}}{\!}\lvert h_{\mathrm{u}}(\mathbf{r})\rvert^{2}\mathrm{d}\mathbf{r}}\right).$ (73)
Proof:
Please refer to Appendix -D for more details. ∎
@@
$\begin{split}\rho_{\mathrm{u}}&=\frac{\pi^{2}L_{x}L_{z}}{4N^{2}}\sum\nolimits_{n=1}^{N}{\sum\nolimits_{n^{\prime}=1}^{N}{\sqrt{\left(1-\xi_{n}^{2}\right)(1-\xi_{n^{\prime}}^{2})}}}\\ &\times\hat{h}_{\mathrm{c}}^{*}\Big{(}\frac{\xi_{n}-1}{2}L_{x},\frac{\xi_{n^{\prime}}}{2}L_{z}\Big{)}\hat{h}_{\mathrm{s}}\Big{(}\frac{\xi_{n}-1}{2}L_{x},\frac{\xi_{n^{\prime}}}{2}L_{z}\Big{)}.\end{split}$ (77)
@@ IV-C1 Performance of Communications
Without loss of generality, we focus on a single slot and assume $\lvert s_{{\rm{s}},\ell}\rvert^{2}=1$ for $\ell=1,\ldots,L$. From a worst-case design perspective, the interference-plus-noise term $z_{{\rm{s}},\ell}=\sqrt{P_{\rm{s}}}s_{{\rm{s}},\ell}\int_{{\mathcal{A}}_{\rm{r}}}v_{\rm{c}}^{*}({\mathbf{r}})a_{\rm{r}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\beta_{\rm{s}}\sqrt{\int_{\mathcal{A}_{\mathrm{t}}}\lvert a_{\mathrm{t}}(\mathbf{t})\rvert^{2}\mathrm{d}\mathbf{t}}+\int_{{\mathcal{A}}_{\rm{r}}}v_{\rm{c}}^{*}({\mathbf{r}})n_{\ell}(\mathbf{r}){\rm{d}}{\mathbf{r}}$ is modeled as Gaussian noise [43], with $z_{{\rm{s}},\ell}\sim{\mathcal{CN}}(0,\sigma_{{\rm{s}},\ell}^{2})$ and
@@
This problem can be solved using a subspace approach similar to that used for problem (68). Consequently, we obtain the uplink CR achieved by the S-C SIC design with the optimal detector, which is expressed as follows:
@@ IV-D Rate-Region Characterization
Therefore, the achievable SR-CR region for the uplink CAPA-based ISAC system is characterized as follows:
@@ V Numerical Results
Numerical results are presented to evaluate the performance of both sensing and communication in CAPA-based downlink and uplink ISAC systems. Unless specified otherwise, the simulation parameters are set as follows: $\lambda=0.125$ m, $L_{x}=L_{z}=0.5$ m ($\lvert{\mathcal{A}}_{\rm{t}}\rvert=\lvert{\mathcal{A}}_{\rm{r}}\rvert=0.25$ m2), $\lvert{\mathcal{A}}_{\rm{r}}\rvert=\frac{\lambda^{2}}{4\pi}$, $\frac{P}{\sigma_{\rm{c}}^{2}}=\frac{P}{\sigma_{\rm{s}}^{2}}=\frac{P_{\rm{c}}}{\sigma^{2}}=\frac{P_{\rm{s}}}{\sigma^{2}}=10$ dB, $\eta=120\pi~{}\Omega$, $L=8$, $\alpha_{\rm{s}}=1$, and $T=100$. The locations of the target and CU are given by $(r_{\rm{s}},\theta_{\rm{s}},\phi_{\rm{s}})=(10~{}\rm{m},\frac{\pi}{4},\frac{\pi}{4})$ and $(r_{\rm{c}},\theta_{\rm{c}},\phi_{\rm{c}})=(20~{}\rm{m},\frac{\pi}{3},\frac{\pi}{3})$, respectively.
@@
  • SPDA-based ISAC: In the SPDA scheme, both the transmit and receive continuous apertures are partitioned into $\lfloor\frac{L_{x}}{d}\rfloor\times\lfloor\frac{L_{z}}{d}\rfloor$ spatially discrete elements, each with a size of $A_{\rm{s}}=\frac{\lambda^{2}}{4\pi}$ [27]. Here, $d=\frac{\lambda}{2}$ is the inter-element spacing. By denoting $d_{\mathrm{s}}=\frac{\sqrt{A_{\mathrm{s}}}}{2}$, the apertures of the transmit and receive arrays are given by
@@
    where $\mathcal{N}_{x}\triangleq\left\{1,...,\lfloor\frac{L_{x}}{d}\rfloor\right\}$ and $\mathcal{N}_{z}\triangleq\left\{1,...,\lfloor\frac{L_{z}}{d}\rfloor\right\}$.
@@
(a) CR versus transmit SNR.
(b) SR versus transmit SNR.
Figure 3: Downlink performance versus transmit SNR.
V-A Downlink Performance
Figure 4(a) and Figure 4(b) illustrate the downlink CR and SR achieved by ISAC under both CAPA and SPDA schemes, which are plotted against the BS’s aperture size. As can be seen from these two graphs, both the C-C CR and the S-C SR increase monotonically with the aperture size. This is because these designs focus energy in optimal directions for communications and sensing, respectively, which results in enhanced array gain and performance as the aperture size grows. In contrast, the C-C SR and S-C CR do not exhibit a strictly monotonic trend. Since the beamformers in these cases are not perfectly aligned with the respective sensing or communication channels, variations in the correlation term $\rho_{\mathrm{d}}=\int_{\mathcal{A}_{\mathrm{t}}}{a_{\mathrm{t}}(\mathbf{t})h_{\mathrm{d}}^{*}(\mathbf{t})}\mathrm{d}\mathbf{t}$ cause fluctuations in the array gain and performance rates. We also find that for each aperture size, CAPA outperforms SPDA, which reinforces the trends observed in Figure 3.

### v2

@@ III-C2 Performance of Sensing
Note that the asymptotic analysis of the S-C design is similar to that of the C-C case, which is omitted here for brevity, while the associate simulation results are presented in Section V.
@@ III-D Pareto-Optimal Design
In practical applications, the beamformer $w({\mathbf{t}})$ can be designed to meet various quality-of-service (QoS) requirements, which creates a trade-off between communication and sensing performance. To evaluate this trade-off, we investigate the Pareto boundary of the achievable SR-CR region in the downlink CAPA-based ISAC system. The Pareto boundary comprises SR-CR pairs where any improvement in one rate can only be achieved by sacrificing performance in the other [51]. This boundary characterizes the optimal trade-off between communication and sensing capabilities, and provides valuable insights for performance optimization.
It is worth noting that the donwlink CR is a monotone increasing function with respect to $\hat{\gamma}_{\rm{c}}({{w}}({\mathbf{t}}))=\lvert\int_{{\mathcal{A}}_{\rm{t}}}h_{\rm{d}}({\mathbf{t}}){{w}}({\mathbf{t}}){\rm{d}}{\mathbf{t}}\rvert^{2}$, while the SR is a monotone increasing function with respect to $\hat{\gamma}_{\rm{s}}({{w}}({\mathbf{t}}))=\lvert\int_{{\mathcal{A}}_{\rm{t}}}a_{\rm{t}}(\mathbf{t})w({\mathbf{t}}){\rm{d}}{\mathbf{t}}\rvert^{2}$. Therefore, the achievable SR-CR region can be determined from the region defined by $\hat{\gamma}_{\rm{c}}({{w}}({\mathbf{t}}))$ and $\hat{\gamma}_{\rm{s}}({{w}}({\mathbf{t}}))$. In particular, any pair located on the Pareto boundary of this region can be identified by solving the following problem:
@@
Please refer to Appendix -C for more details. ∎
@@
Although problem (58) is more mathematically tractable than problem (54), it remains non-convex. To solve this, we derive a closed-form solution using the Karush–Kuhn–Tucker (KKT) conditions.
@@
Please refer to Appendix -D for more details. ∎
@@
Note that problem (54) can alternatively be solved using the Fourier-based approach [30, 31, 32, 38], where the spatial-domain CAPA channel is transformed into its wavenumber-domain representation, converting the original continuous problem into a discretized one. However, such discretization introduces inherent approximation loss and scales poorly with CAPA aperture size and carrier frequency, resulting in high computational complexity. Specifically, obtaining the Fourier coefficients for both sensing and communication channels requires calculating $2\times\left(2\lceil\frac{L_{x}}{\lambda}\rceil+1\right)\times\left(2\lceil\frac{L_{z}}{\lambda}\rceil+1\right)$ integrals. In contrast, our subspace-based formulation directly optimizes the continuous source current pattern, which requires calculating only one integral for the channel correlation $\rho_{\rm{d}}$, significantly reducing complexity. Moreover, it provides a closed-form solution that better captures the system’s continuous nature and offers clearer insight into performance limits.
@@ IV-B1 Performance of Sensing
We assume that the following aggregate interference-plus-noise term:
@@
is Gaussian noise333 We assume $z_{c,l}$ as Gaussian noise for two reasons. First, the transmitted data symbol $s_{{\rm{c}},\ell}$ is often considered as complex Gaussian random variables with zero mean and unit variance, i.e., $s_{{\rm{c}},\ell}\sim\mathcal{CN}(0,1)$ [53, 54]. This assumption is widely adopted as it reflects the capacity-achieving input distribution under power constraints [55]. Also, modeling the transmit signal as Gaussian is a valid approximation when many modulated symbols are superimposed (e.g., in large-scale antenna arrays) due to the central limit theorem [56]. The second reason is that the Gaussian distribution maximizes the entropy among all distributions with the same power, and thus minimizes the mutual information between the input and output of the channel [55, 57]. Therefore, modeling the aggregated interference-plus-noise as Gaussian represents the worst-case noise, which provides a conservative and analytically tractable lower bound of the system performance. , with $z_{{\rm{c}},\ell}\sim{\mathcal{CN}}(0,\sigma_{{\rm{c}},\ell}^{2})$ and
@@
$a_{\rm{r}}(\mathbf{r})$ $=a_{\rm{r},1}\psi_{1}(\mathbf{r})+a_{\rm{r},2}\psi_{2}(\mathbf{r}).$ (69c)
@@
where ${\mathbf{v}}=[v_{1},v_{2}]^{\mathsf{T}}$, ${\mathbf{h}}_{\rm{u}}=[h_{\rm{u},1},h_{\rm{u},2}]^{\mathsf{T}}$, and ${\mathbf{a}}_{\rm{r}}=[a_{\rm{r},1},a_{\rm{r},2}]^{\mathsf{T}}$. This is now a classical matrix-form maximization of a generalized Rayleigh quotient, whose optimal solution is given by
@@
By applying the Woodbury matrix identity to this solution and based on (69), we obtain the optimal detector $v_{\rm{s}}^{\star}({\mathbf{r}})$ as follows.
@@
In the C-C SIC, the optimal detector for the sensing signal is given by
$v_{\mathrm{s}}^{\star}(\mathbf{r})\propto a_{\mathrm{r}}(\mathbf{r})-\frac{\tilde{\gamma}_{\mathrm{c}}\int_{\mathcal{A}_{\mathrm{r}}}{a_{\mathrm{r}}(\mathbf{r})h_{\mathrm{u}}^{*}(\mathbf{r})}\mathrm{d}\mathbf{r}}{1+\tilde{\gamma}_{\mathrm{c}}\int_{\mathcal{A}_{\mathrm{r}}}{\left|h_{\mathrm{u}}(\mathbf{r})\right|^{2}}\mathrm{d}\mathbf{r}}h_{\mathrm{u}}(\mathbf{r}).$ (72)
@@
$\begin{split}\rho_{\mathrm{u}}&=\frac{\pi^{2}L_{x}L_{z}}{4N^{2}}\sum\nolimits_{n=1}^{N}{\sum\nolimits_{n^{\prime}=1}^{N}{\sqrt{\left(1-\xi_{n}^{2}\right)(1-\xi_{n^{\prime}}^{2})}}}\\ &\times\tilde{h}_{\mathrm{c}}^{*}\Big{(}\frac{\xi_{n}-1}{2}L_{x},\frac{\xi_{n^{\prime}}}{2}L_{z}\Big{)}\tilde{h}_{\mathrm{s}}\Big{(}\frac{\xi_{n}-1}{2}L_{x},\frac{\xi_{n^{\prime}}}{2}L_{z}\Big{)}.\end{split}$ (76)
@@ IV-C1 Performance of Communications
Without loss of generality, we focus on a single slot and assume $\lvert s_{{\rm{s}},\ell}\rvert^{2}=1$ for $\ell=1,\ldots,L$. The interference-plus-noise term $z_{{\rm{s}},\ell}=\sqrt{P_{\rm{s}}}s_{{\rm{s}},\ell}\int_{{\mathcal{A}}_{\rm{r}}}v_{\rm{c}}^{*}({\mathbf{r}})a_{\rm{r}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\beta_{\rm{s}}\sqrt{\int_{\mathcal{A}_{\mathrm{t}}}\lvert a_{\mathrm{t}}(\mathbf{t})\rvert^{2}\mathrm{d}\mathbf{t}}+\int_{{\mathcal{A}}_{\rm{r}}}v_{\rm{c}}^{*}({\mathbf{r}})n_{\ell}(\mathbf{r}){\rm{d}}{\mathbf{r}}$ can be regarded as Gaussian noise, with $z_{{\rm{s}},\ell}\sim{\mathcal{CN}}(0,\sigma_{{\rm{s}},\ell}^{2})$ and
@@
This problem can be solved using a subspace approach similar to that used for problem (67).
Lemma 4.
The optimal detector $v_{\mathrm{c}}^{\star}(\mathbf{r})$ is given by
$\!\!v_{\mathrm{c}}^{\star}(\mathbf{r})\!\propto\!h_{\mathrm{u}}(\mathbf{r})\!-\!\frac{\tilde{\gamma}_{\mathrm{s}}\int_{\mathcal{A}_{\mathrm{t}}}\!{\left|a_{\mathrm{t}}(\mathbf{t})\right|^{2}}\mathrm{d}\mathbf{t}\!\int_{\mathcal{A}_{\mathrm{r}}}\!{a_{\mathrm{r}}^{*}(\mathbf{r})h_{\mathrm{u}}(\mathbf{r})}\mathrm{d}\mathbf{r}}{1\!+\!\tilde{\gamma}_{\mathrm{s}}\int_{\mathcal{A}_{\mathrm{t}}}\!{\left|a_{\mathrm{t}}(\mathbf{t})\right|^{2}}\mathrm{d}\mathbf{t}\!\int_{\mathcal{A}_{\mathrm{r}}}\!{\left|a_{\mathrm{r}}(\mathbf{r})\right|^{2}}\mathrm{d}\mathbf{r}}a_{\mathrm{r}}(\mathbf{r}).$ (82)
Consequently, the uplink CR achieved by the S-C SIC design can be expressed as follows:
@@ IV-D Rate-Region Characterization
Consequently, the achievable SR-CR region for the uplink CAPA-based ISAC system can be characterized as follows:
@@ V Numerical Results
Numerical results are presented to evaluate the performance of both sensing and communication in CAPA-based downlink and uplink ISAC systems. Unless specified otherwise, the simulation parameters are set as follows: $\lambda=0.125$ m, $L_{x}=L_{z}=0.5$ m ($\lvert{\mathcal{A}}_{\rm{t}}\rvert=\lvert{\mathcal{A}}_{\rm{r}}\rvert=0.25$ m2), $\lvert{\mathcal{A}}\rvert=\frac{\lambda^{2}}{4\pi}$, $\frac{P}{\sigma_{\rm{c}}^{2}}=\frac{P}{\sigma_{\rm{s}}^{2}}=\frac{P_{\rm{c}}}{\sigma^{2}}=\frac{P_{\rm{s}}}{\sigma^{2}}=10$ dB, $\eta=120\pi~{}\Omega$, $L=8$, $\alpha_{\rm{s}}=1$, and $T=50$. The locations of the target and CU are given by $(r_{\rm{s}},\theta_{\rm{s}},\phi_{\rm{s}})=(10~{}\rm{m},\frac{\pi}{4},\frac{\pi}{4})$ and $(r_{\rm{c}},\theta_{\rm{c}},\phi_{\rm{c}})=(20~{}\rm{m},\frac{\pi}{3},\frac{\pi}{3})$, respectively.
(a) CR versus transmit SNR.
(b) SR versus transmit SNR.
Figure 3: Downlink performance versus transmit SNR.
@@
  • SPDA-based ISAC: In the SPDA scheme, both the transmit and receive continuous apertures are partitioned into $\lfloor\frac{L_{x}}{d}\rfloor\times\lfloor\frac{L_{z}}{d}\rfloor$ spatially discrete elements, each with a size of $A_{\rm{s}}=\frac{\lambda^{2}}{4\pi}$. Here, $d=\frac{\lambda}{2}$ is the inter-element spacing. By denoting $d_{\mathrm{s}}=\frac{\sqrt{A_{\mathrm{s}}}}{2}$, the apertures of the transmit and receive arrays are given by
@@
    where $\mathcal{N}_{x}\triangleq\left\{1,...,\lfloor\frac{L_{x}}{d}\rfloor\right\}$ and $\mathcal{N}_{z}\triangleq\left\{1,...,\lfloor\frac{L_{z}}{d}\rfloor\right\}$. We assume that the SPDA is based on a fully digital architecture, where each antenna element is connected with a dedicated RF chain. On the other hand, the CAPA considered in our work only requires a single RF chain, as the number of spatially multiplexed signals is one.
@@
V-A Downlink Performance
Figure 4(a) and Figure 4(b) illustrate the downlink CR and SR achieved by ISAC under both CAPA and SPDA schemes, which are plotted against the BS’s aperture size. As can be seen from these two graphs, both the C-C CR and the S-C SR increase monotonically with the aperture size. This is because these designs focus energy in optimal directions for communications and sensing, respectively, which results in enhanced array gain and performance as the aperture size grows. In contrast, the C-C SR and S-C CR do not exhibit a strictly monotonic trend. Since the beamformers in these cases are not perfectly aligned with the respective sensing or communication channels, variations in the correlation term $\rho_{\mathrm{d}}=\int_{\mathcal{A}_{\mathrm{t}}}{a_{\mathrm{t}}(\mathbf{t})h_{\mathrm{d}}^{*}(\mathbf{t})}\mathrm{d}\mathbf{t}$ cause fluctuations in the array gain and performance rates. Moreover, as the aperture size grows, both the CRs ans SRs eventually converge to finite values larger than zero, adhering to the principle of energy conservation, which is consistent with the statements in Remark 2.
@@ V-A Downlink Performance
While the Fourier-based approach can achieve a Pareto boundary comparable to the proposed subspace method, Table I presents the runtime for each method to obtain the Pareto boundary under varying CAPA aperture sizes and carrier frequencies. All experiments were conducted on a Windows 10 laptop equipped with an AMD Ryzen 7 5800H processor, using MATLAB R2023a. Each method was executed 100 times, and the average runtime was recorded. We observe from the table that the runtime of the Fourier-based approach increases significantly with the aperture size and frequency. This is attributed to the rapid growth in the number of integrals—computed via Gaussian quadrature—required by the Fourier method as these parameters scale. For example, when increasing the aperture size and frequency from $0.25\,\text{m}^{2}$ and $2.4$ GHz to $1\,\text{m}^{2}$ and $10$ GHz, the number of integral calculated within the Fourier-based approach rises from $162$ to $9522$. In contrast, the subspace approach consistently maintains a low runtime of approximately $0.005$ seconds, as it involves only a single integral whose complexity is independent of both aperture size and frequency. These results underscore the computational efficiency of the proposed subspace approach.
TABLE I: Comparison of runtime for different methods
Frequency $L_{x}=L_{z}=0.5\,\text{m}$ $L_{x}=L_{z}=0.7\,\text{m}$ $L_{x}=L_{z}=1\,\text{m}$
Subspace Fourier Subspace Fourier Subspace Fourier
2.4 GHz 0.005 s 0.163 s 0.005 s 0.335 s 0.005 s 0.642 s
5 GHz 0.005 s 0.721 s 0.005 s 1.272 s 0.005 s 2.495 s
10 GHz 0.005 s 2.539 s 0.005 s 4.926 s 0.005 s 9.885 s